Circular No. 7337

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999go IN NGC 1376 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in NGC 1376. SN 1999go was discovered and confirmed on unfiltered images taken on Dec. 23.2 and 24.2 UT, both showing the new object at mag ∼ 15.5. SN 1999go is located at α = 3^{h}37^{m}05^{s}.72, δ = −5^{o}02′19′′.0 (equinox 2000.0), which is 2′′.6 west and 14′′.8 north of the nucleus of NGC 1376. A KAIT image of the same field on Dec. 15.2 showed nothing at the position of the supernova (limiting mag ∼ 18.5).

SUPERNOVA 1999gp IN UGC 1993 Papenkova and Li also report the discovery by LOSS of an apparent

supernova (mag ∼ 18.0) on Dec. 23.2 UT (and confirmed on Dec. 24.2 at mag ∼ 17.2). SN 1999gp is located at α = 2^{h}31^{m}39^{s}.15, δ = +39^{o}22′52′′.3 (equinox 2000.0), which is 11′′.4 west and 9′′.9 north of the nucleus of UGC 1993. A KAIT image of the same field on Dec. 18.2 showed nothing at the position of the supernova (limiting mag ∼ 19.0).

SUPERNOVA 1999gq IN NGC 4523 Papenkova and Li also report the discovery by LOSS of an apparent

supernova (mag ∼ 14.5) on Dec. 23.6 UT (confirmed on Dec. 24.6 at mag ∼ 14.6). SN 1999gq is located at α = 12^{h}33^{m}48^{s}.32, δ = +15^{o}10′48′′.2 (equinox 2000.0), which is ∼ 5′′ east and 46′′ north of the diffuse nucleus of NGC 4523. A KAIT image of the same field on Apr. 10.3 showed nothing at the position of the supernova (limiting mag ∼ 19.0).

V1494 AQUILAE Visual magnitude estimates: Dec. 12.054 UT, 6.3 (R. Keen, Mount

Thorodin, CO); 13.719, 6.2 (S. Baroni, Milan, Italy); 15.925, 6.4 (R. Lourencon, Jundiai, Brazil); 18.658, 6.7 (B. H. Granslo, Oslo, Norway); 19.86, 7.1 (J. Carvajal, Madrid, Spain); 21.701, 7.3 (S. Foglia, Milan, Italy); 23.703, 7.2 (K. Hornoch, Lelekovice, Czech Republic).

1999 December 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green